

********************************************************************
UARS WEEKLY STATUS REPORT - 19 March 2004
********************************************************************
***********
SOLSTICE
***********
During this week SOLSTICE continued to operate in good health gathering
solar data on all available orbital tracking periods. On average, we
received 49% of each day's data amounting to about 522 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg II,
occultation, solar pointing offset and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
Congratulations and thanks to those operations folks who have made
possible the 3-TDRS operation. There is a substantial improvement!
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 06 March 2004 (Orbit 68308, GMT Day 066) through 12
March 2004 (Orbit 68412, GMT Day 072).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 068/1403 68346 PEM AXIS ON
PEM 068/1433 68346 PEM ZEPS & NEPS FULL TIME
The beta angle ranged between the angles of 52.1 degrees to 60.2
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 29 March 2004 (DOY 089).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.26 degrees C. Additional
battery performance data for each battery on 05 March and 12 March
2004 are provided in Appendix D.
The clock error ranged from 11.1 to -10.5 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
PEM = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT completed testing the new version of the CMS code this week.
All CMS processes were examined and functioned properly. The UARS
mini-ORR was held on 09 March 2004 to review the results of the
testing and outline the plan to utilize the new 3-TDRS functionality.
The review board gave approval to proceed with the full 3-TDRS
capability. The FOT will transition the CMS to the new software on 15
March 2004 and then begin three contacts per orbit after 1400 Z on 17
March 2004.
MEETINGS
The FOT held the UARS mini-ORR on 09 March 2004.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 March 2004 (DOY 089). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 March 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68346 068/1433 04-075 PEM PEM ZEPS & NEPS TO FULL TIME
68346 068/1403 04-076 PEM PEM AXIS ON
68347 068/1615 04-077 PEM PMON / PEM
68354 069/0235 04-078 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68395 071/2010 04-079 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 March 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 March 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68354 069/0237 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68395 071/2015 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 March 2004
05 March 2004 - GMT Day 065 Beta = 50.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +16.8/-100.8 +5.14/+2.91 +25.2/-14.8 25.8 1.09 17.2
3 +11.2/-095.2 +3.87/+1.97 +24.8/-15.2 25.6 1.08 17.0
12 March 2004 - GMT Day 072 Beta = 60.2 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-022.4 +3.87/+2.60 +22.0/-15.2 26.1 1.02 17.8
3 +11.2/-033.6 +2.91/+1.65 +21.2/-15.2 25.9 1.02 17.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None This Report Period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 March 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-014 FOT See List Mar 04 30832 Generic Late Acquisition